OpenStudy (anonymous):

Can somebody tell me how the Taylor expansion of\[\ln\left(\frac{x}{n}+1\right)=\frac{x}{n}-\frac{x^2}{2n^2}+\frac{x^3}{3n^3}-\cdots?\] According to the definition,\[f(a)+\frac{f'(a)}{1!}(x-a)+\frac{f''(a)}{2!}(x-a)^2+\cdots,\]but I can't seem to make the connection.

OpenStudy (anonymous):

i can come back to this but if you want an immediate guess i would say take the derivative, probably get an easy expansion, and then integrate term by term that is just a guess though

OpenStudy (anonymous):

ok i am confused why does your function have an "n" in it??

OpenStudy (anonymous):

damn it is not working. i thought you could integrate this thing term by term and get what you want, but maybe i made a calculation error somewhere

OpenStudy (anonymous):

we could try the expansion as you wrote it we are expanding about 0 so \[f(0)=\ln(1)=0\] first term is 0 \[f'(0)=\frac{n}{n+0}=1\] so second term is x

OpenStudy (anonymous):

something is screwy here

OpenStudy (anonymous):

are you sure there in an "n" in this?

OpenStudy (anonymous):

no it is right, i am screwy. hold on and let me see what we can do

OpenStudy (anonymous):

lord it would help if i knew how to differentiate sorry

OpenStudy (anonymous):

\[f(x)=\ln(\frac{x}{n}+1)\] \[f'(x)=\frac{1}{x+n}\] i forgot the damned chain rule!!

OpenStudy (anonymous):

\[f'(0)=\frac{1}{n}\] so the expansion starts with \[\frac{x}{n}\]

OpenStudy (anonymous):

\[f''(x)=-\frac{1}{(x+n)^2}\] \[f''(0)=-\frac{1}{n^2}\] so second term of the expansion is \[-\frac{x^2}{2n^2}\]

OpenStudy (anonymous):

you got this or should i keep going?

OpenStudy (anonymous):

beautiful. it all makes sense now. thank you so very much!!